Consider the line x + 6y= -3.
6y=-x-3
y=-1/6x-1/2
What is the slope of a line perpendicular to this line?
Opposite of -1/6= 6!
What is the slope of a line parallel to this line?
It's just the same as the original, -1/6.
I hope I helped!

Answer by Miceandwine(15)

Consider the line x + 6y= -3.
6y=-x-3
y=-1/6x-1/2
What is the slope of a line perpendicular to this line?
Opposite of -1/6= 6!
What is the slope of a line parallel to this line?
It's just the same as the original, -1/6.
I hope I helped!